OpenStudy (anonymous):

If f(x) varies directly with x, and f(x) = 56 when x = 8, find the value of f(x) when x = 2

OpenStudy (anonymous):

marry me yes? :DDD

OpenStudy (anonymous):

uh maybe if you answer the question correctly

OpenStudy (anonymous):

Alright so if f(x) = 58 when x = 8 we can conclude that if we divide 58 by 8 we can find the f. In this case f = 7. So the question now is what is 7(2)?

OpenStudy (anonymous):

Woah woah woah, where did we get 58 from....

OpenStudy (anonymous):

My brain thought it said 58. My derp. Give me a second to correct my mistake.

OpenStudy (anonymous):

Alright F is still 7. Our equation is still 7(2). I just hit the wrong key. The woes of a new key bored. Can you solve 7(2)?

OpenStudy (anonymous):

56/8= 7 since 8*7=56 so f(x) would be 7(2) and 7*2=14...so im guessing its 14

OpenStudy (anonymous):

^^ Thats the answer xD

OpenStudy (anonymous):

Lol I was bout to answer w/ 14 but thanks anyway I appreciate it!

OpenStudy (anonymous):

your welcome @Chelleyhi